OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations: 4y = x − 13 2x + 16y = 20 (14, 1/4) (-3, 1) (12, -1/4) (5, -2)

OpenStudy (anonymous):

@pengembara_bumi1 help please

OpenStudy (anonymous):

-x+4y=-13 2x+16y=20 Multiply the top equation by 2 to cancel out the x. -2+8y=-26 2x+16y=20 Cancel the x and add the ys and constants 24y=-6 Divide. y=6/24 Simplify y=1/4 Take the y value and plug it in for x in one of the original equations.

OpenStudy (anonymous):

sorry, plug it in for y in an original

OpenStudy (anonymous):

thanks can you help me on the next one